Identify the vertext for graph of y=4(x+3)^2+4
Found 2 solutions by scott8148, lwsshak3:
Answer by scott8148(6628)

the general vertex form is ___ y = a(x - h)^2 + k ___ (h,k) is the vertex

Answer by lwsshak3(11628)

Identify the vertex for graph of y=4(x+3)^2+4
**
Standard form of equation for a parabola that opens upwards: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of the vertex. A is a multiplier which affects the slope or steepness of the curve.
For given equation, y=4(x+3)^2+4
vertex: (-3,4)